### v5

Abstract
Defining pathologies automatically from medical images aids the understanding of the emergence and progression of diseases, and such an ability is crucial in clinical diagnostics. However, existing deep learning models heavily rely on expert annotations and lack generalization capabilities in open clinical environments. In this study, we present a generalizable vision-language model for Annotation-Free pathology Localization (AFLoc). The core strength of AFLoc lies in its extensive multi-level semantic structure-based contrastive learning, which comprehensively aligns multi-granularity medical concepts from reports with abundant image features, to adapt to the diverse expressions of pathologies and unseen pathologies without the reliance on image annotations from experts. We conducted primary experiments on a dataset of 220K pairs of image-report chest X-ray images, and performed extensive validation across six external datasets encompassing 20 types of chest pathologies. The results demonstrate that AFLoc outperforms state-of-the-art methods in both annotation-free localization and classification tasks. Additionally, we assessed the generalizability of AFLoc on other modalities, including histopathology and retinal fundus images. Extensive experiments show that AFLoc exhibits robust generalization capabilities, even surpassing human benchmarks in localizing five different types of pathological images. These results highlight the potential of AFLoc in reducing annotation requirements and its applicability in complex clinical environments.
keywords:
Annotation-Free Deep Learning, pathology localization, Vision-Language Pre-Training
Accurate diagnosis and precise pathology localization in medical images are paramount for optimizing treatment effectiveness and interventions. These technologies are instrumental in personalized medicine, facilitating customized treatment approaches that improve patient outcomes and mitigate the possibility of diagnostic errors. By pinpointing the exact location and extent of abnormalities, clinicians can make informed decisions that lead to more targeted therapies and improved prognoses for patients [1, 2, 3, 4].
Over the past decade, supervised deep learning methods have significantly accelerated advancements in disease localization [5, 6, 7, 8]. However, the efficacy of these methods heavily relies on extensively annotated training datasets, which require domain experts to invest considerable time [9]. Specifically, clinical localization tasks often require experienced clinicians to meticulously annotate numerous precise bounding boxes or perform pixel-wise delineations of localized pathology areas. This annotation process is costly, particularly in resource-constrained clinical settings, and algorithms frequently struggle to generalize to diverse datasets.
Several methods have been proposed to reduce the reliance on large annotated datasets [10, 11, 12, 13]. Initially, these methodologies acquire general visual representations through self-supervised learning from image datasets, followed by fine-tuning on smaller annotated datasets. This approach enables models to achieve high performance on specific tasks while significantly decreasing the need and cost of data labeling [13]. Moreover, saliency-based methods [14, 15, 16, 17] have been developed to reduce annotation costs in pathology localization tasks by allowing coarse localization of target categories in models trained with image-level annotations. However, these methods still require annotations for specific downstream tasks. This requirement is particularly challenging in flexible and dynamic clinical environments, especially for newly emerging diseases (e.g., COVID-19), where deployed models may fail to perform effectively [18, 19, 20].
In recent years, unsupervised deep learning methods have gained increasing attention due to their independence from annotated datasets, particularly in the field of anomaly detection [21, 22, 23, 24]. These methods typically train models using only healthy samples, learning the distribution of normal anatomical structures, which enables the identification of abnormal pathology samples during the testing phase [21, 22]. They are particularly effective for data with simple structures and low inter-sample variance, allowing them to learn normative distributions and achieve excellent anomaly detection performance [23, 24, 25]. However, challenges such as the high heterogeneity of pathology images, significant similarities between different pathologies, and large variations in contrast for the same lesions reduce the usability of these methods in complex scenarios, thereby hindering their practical application in real medical environments [25].
A promising approach is the development of medical visual-language pre-training methods [26, 27, 28, 29, 19, 30, 31]. These methods establish effective correlations between medical reports and medical images, allowing them to flexibly localize disease types not encountered during pre-training without requiring additional customized annotations [19]. However, achieving precise pathology localization solely through the combination of medical images and clinical reports remains challenging. A primary obstacle is the lack of explicit pathology localization markers in clinical reports, which often provide only coarse information such as ‘upper’ or ‘left’ to indicate disease location. Moreover, clinical descriptions by clinicians are subjective and variable, further complicating the task of accurately extracting and localizing diseases in medical images. To address this challenge, several methods have been proposed to integrate finer-grained information. For instance, GLoRIA [30] extracts the correlation of the image’s regions and paired words in reports to learn global and local representations of images. MedKLIP [19] utilizes well-defined medical vocabulary knowledge bases to provide supervision at the entity level through triplet training paradigms. However, these fine-grained methods typically focus on individual levels of medical concepts and may overlook the variable meanings of concepts in different contexts. Therefore, these approaches may struggle to adapt to the diverse expressions of disease descriptors in clinical practice, often requiring customized textual cues to enhance localization performance.
In this study, we propose AFLoc, a visual-language model based on contrastive learning aimed at alleviating the need for costly pathology localization annotations. AFLoc can autonomously perform pathology localization and clinical diagnosis with medical images. Unlike traditional global semantic alignment strategies [29, 28], AFLoc introduces a contrastive learning framework with a multi-level semantic alignment component, facilitating the comprehensive alignment of medical concepts from reports with image features. Specifically, the image encoder generates three levels of features: shallow local features, deep local features, and global features, which are aligned with word-level, sentence-level, and report-level features extracted by the text encoder. We extensively validated AFLoc across three types of medical image datasets, including Chest X-Ray (6 external datasets), histopathology (3 external datasets), and retinal fundus images. Our results show that AFLoc outperforms state-of-the-art methods in localization and clinical diagnostic tasks across different modalities. We hope that this study can help address the challenges posed by annotation scarcity and modality diversity in clinical environments, while providing new insights for the design of future clinical open-environment methods.
a
b
Results
We primarily evaluated AFLoc on two tasks: pathology localization and clinical diagnosis. Three modalities were considered, including Chest X-ray, histopathology, and retinal fundus images. In addition, we conducted extensive ablation studies to evaluate the effects of different prompts, text granularities, and text encoders on AFLoc’s performance. These experiments are discussed in detail in subsequent sections.
Annotation-Free Localization Across Three Medical Modalities
AFLoc is capable of performing zero-shot localization tasks without requiring any annotations. In this section, we present AFLoc’s annotation-free localization performance across different clinical modalities. Each modality has been pre-trained using its corresponding image-report datasets, ensuring modality-specific knowledge is effectively captured and utilized. Here, we consider that ’pathology localization’ encompasses both heatmap predictions and binary mask results, as some clinical applications may benefit more from the visualization provided by heatmaps rather than discrete segmentation. We quantitatively evaluate each predicted heatmap using Contrast-to-Noise Ratio (CNR) [28] and assess the thresholded segmentation masks using Intersection over Union (IoU) [32] and Dice similarity coefficient metrics. In the following, we will present the results of the three modalities separately.
Localization Performance on Chest X-Ray In clinical practice for chest diseases, doctors often need to perform detailed assessments of Chest X-rays for early screening and surgical planning. For instance, airspace opacity can indicate lung infections, inflammation, or tumors [33]. Atelectasis leads to restricted gas exchange, causing hypoxemia and severely affecting respiratory function [34]. However, since these diseases often lack clear boundaries and have contrast levels similar to surrounding tissues, doctors typically need to invest significant effort in distinguishing them. Therefore, automated diagnostic systems play a crucial role in aiding diagnosis.
We evaluated the localization performance of AFLoc on Chest X-rays using four external datasets: RSNA Pneumonia [35], MS-CXR [28], CheXlocalize [14], and COVID Rural [36, 37]. These datasets cover 13 common thoracic pathologies: pneumonia, airspace opacity, atelectasis, cardiomegaly, consolidation, edema, enlarged cardiomediastinum, lung lesion, pleural effusion, pneumothorax, lung opacity, support devices, and COVID-19.
We first elaborate AFLoc’s localization performance on the RSNA Pneumonia dataset (Fig. 2, Extended Data Table 3). AFLoc achieves better results than all comparative methods across all three evaluation metrics. Specifically, compared to the respective best-performing comparative methods, AFLoc boosts the IoU by 7.9% from 0.317 (95% CI: 0.309, 0.325) to 0.342 (95% CI: 0.332, 0.350), enhances the Dice coefficient by 4.1% from 0.465 (95% CI: 0.455, 0.474) to 0.484 (95% CI: 0.474, 0.495), and elevates CNR by 18.3% from 1.300 (95% CI: 1.262, 1.339) to 1.538 (95% CI: 1.496, 1.580). These results highlight AFLoc’s superior ability to localize pathological regions.
Similarly, the results on the MS-CXR dataset (Fig. 2, Extended Data Table 6) show that our AFLoc outperforms the existing vision-language pre-training approaches (GLoRIA, BioViL, and MedKLIP) by a clear margin. For example, AFLoc improves the IoU (0.324, 95% CI: 0.298, 0.350) by more than 6% compared to GLoRIA (0.268, 95% CI: 0.242, 0.295), BioViL (0.228, 95% CI: 0.201, 0.255), and MedKLIP (0.264, 95% CI: 0.240, 0.288). In addition, AFLoc shows consistent performance enhancement across the majority of pathologies (six out of eight) (Fig. 3), confirming its superior pathology localization capability. We also compared AFLoc with saliency-based methods (Grad-CAM, Grad-CAM++, Eigen-CAM, and the human benchmark) on this dataset using published results from CheXlocalize (Fig. 3b). AFLoc outperforms these methods in most cases, achieving a 4% higher average IoU than Grad-CAM (0.318 vs. 0.282). AFLoc also surpasses the human benchmark in localizing five pathologies, including airspace opacity and pleural effusion, demonstrating its potential in medical image analysis.
Finally, we demonstrate AFLoc’s ability to generalize to unseen diseases by using the COVID Rural dataset (X-rays of COVID-19 pneumonia). AFLoc achieves an IoU of 0.211 (95% CI: 0.185, 0.236), surpassing the performance of MedKLIP (0.174, 95% CI: 0.156, 0.192). AFLoc also achieves higher Dice coefficient (0.323, 95% CI: 0.289, 0.355) and CNR (1.062, 95% CI: 0.929, 1.173) when compared to all comparative models. These results underscore AFLoc’s potential for real-world applications of AFLoc in clinical scenarios with emerging conditions.
Localization Performance on Retinal Fundus Images To further demonstrate the generalizability of our AFLoc, we applied it to retinal fundus images. We curated a retinal fundus image dataset containing three prevalent retinal pathologies: choroidal neovascularization (CNV), drusen, and intraretinal hemorrhages. Detecting choroidal neovascularization is essential for the early diagnosis of wet age-related macular degeneration (AMD) [38]. Drusen serves as an indicator of dry AMD [38]. Intraretinal hemorrhages can reflect severe conditions such as diabetic retinopathy or hypertensive retinopathy [39]. Early detection and identification of these pathologies can significantly slow disease progression and facilitate proactive interventions.
The localization results of different methods on these three retinal pathologies are depicted in Fig. 2 (column 4) and Extended Data Table 4. Notably, AFLoc attains the highest IoU, Dice, and CNR scores across the different pathologies. For choroidal neovascularization, AFLoc achieves an IoU of 0.321 (95% CI: 0.303, 0.337), which is much higher than GLoRIA (0.143, 95% CI: 0.130, 0.156) and other comparative methods. Regarding drusen, AFLoc achieves an IoU of 0.147 (95% CI: 0.125, 0.170) and a Dice score of 0.225 (95% CI: 0.192, 0.257), showcasing its proficiency in localizing subtle features. While AFLoc achieves superior scores compared to all comparative methods, its capability to localize intraretinal hemorrhages is relatively weaker than for the other two pathologies, with an IoU of 0.097 (95% CI: 0.091, 0.102). This suggests the necessity for further enhancements in localizing diffuse and less distinct lesions. In Fig. 4, the visualization of retinal localization is presented, with deeper red areas denoting model-predicted disease locations, while the black box signifies clinician annotations. AFLoc accurately localizes concentrated pathologies like choroidal neovascularization and drusen.
In summary, the localization results in this section underscore that AFLoc proves effective not only in chest X-ray images, as discussed in the previous section, but also in retinal fundus images. This showcases its potential as a versatile localization tool capable of generalizing across various imaging modalities.
[…]
Here, we utilized the Quilt-1M dataset [41] to train the AFLoc model for histopathology and tested its performance on the SICAPv2 test set (n=2,122) under the annotation-free localization setting. We compared AFLoc with several state-of-the-art methods, including CLIP [42], BiomedCLIP [43], and three foundation models specialized for histopathology: PLIP [44], QuiltNet [41], and CONCH [45]. Specifically, we applied GradCAM [15] for the comparative methods, as it has been shown to enhance their performance.
As shown in Fig. 2 and Extended Data Table 5, AFLoc achieves the best localization performance across all three evaluation metrics. Specifically, compared to the best-performing comparative method, PLIP, AFLoc demonstrates notable improvements in IoU (0.282, 95% CI: 0.264, 0.300 vs. 0.249, 95% CI: 0.230, 0.269), Dice score (0.396, 95% CI: 0.375, 0.418 vs. 0.352, 95% CI: 0.327, 0.376), and CNR (0.639, 95% CI: 0.572, 0.706 vs. 0.321, 95% CI: 0.249, 0.398). These results suggest that AFLoc is capable of more accurately localizing abnormal tissues compared to both general-purpose and histopathology-specific models.
While the results are encouraging, it is worth noting that histopathological localization remains a highly challenging task due to the complexity and variability of tissue morphology. Nevertheless, AFLoc’s superior performance across all metrics signifies its potential as a robust and generalizable tool for annotation-free pathology localization in histopathology.
Ablation Study of Two Types of Prompts
In the previous section, we validated AFLoc’s ability for annotation-free localization across three imaging modalities using tailored prompts. In clinical settings, there are two prompt generation methodologies: 1) automated prompts based on general-purpose schemes for different imaging modalities, and 2) manually crafted prompts for enhanced precision in localization, as provided by the MS-CXR dataset. This raises the question: to what extent can customized prompts improve AFLoc’s performance?
To explore this, we compared the results obtained using simple and precise prompts on the MS-CXR dataset. For example, the simple prompt for pneumonia can be “findings suggesting pneumonia”, while the precise description can be “severe bibasilar consolidation” or “air space opacity in a right infrahilar location”, reflecting clinical terminology.
Results in Table 1 show that when utilizing precise prompts, all models, including our AFLoc and the two comparative methods (BioViL and GLoRIA), demonstrate enhanced localization performance. For our AFLoc, the utilization of precise prompts leads to a 12.1% increase in IoU (from 0.289 to 0.324), a 10.5% increase in Dice (from 0.418 to 0.462), and a 21.1% increase in CNR (from 1.351 to 1.636). These improvements highlight the advantages of detailed, clinically relevant prompts to enhance localization accuracy. Nevertheless, AFLoc outperforms the two comparative methods regardless of the prompt generation methods.
Fig. 5 illustrates that with simple prompts, models tend to have broader, less accurate focal areas, frequently failing to align with radiologist-annotated regions. In contrast, with precise prompts, AFLoc exhibits enhanced focus, closely matching the annotations made by radiologists. For example, in the case of left lower lobe pneumonia, AFLoc accurately highlights the affected region with precise prompts. Similarly, for the case involving tiny right basilar pneumothorax, much better localization performance is observed when precise prompts are utilized compared to simple prompts.
| Description | BioViL | GLoRIA | AFLoc |
| IoU | |||
| Simple description | 0.187 (0.162,0.213) | 0.240 (0.214,0.265) | 0.289 (0.262,0.314) |
| Precise description | 0.228 (0.201,0.255) | 0.268 (0.242,0.295) | 0.324 (0.298,0.350) |
| Dice | |||
| Simple description | 0.287 (0.252,0.321) | 0.357 (0.322,0.392) | 0.418 (0.383,0.450) |
| Precise description | 0.342 (0.308,0.377) | 0.392 (0.357,0.427) | 0.462 (0.431,0.492) |
| CNR | |||
| Simple description | 0.825 (0.694,0.948) | 1.128 (1.003,1.254) | 1.351 (1.219,1.481) |
| Precise description | 1.083 (0.949,1.219) | 1.287 (1.149,1.421) | 1.636 (1.501,1.772) |
Annotation-Free Diagnosis Across Three Medical Modalities
In this section, we demonstrate AFLoc’s annotation-free diagnostic performance through zero-shot classification, using CXR, retinal fundus images, and histopathology datasets. For comparison, we also present the results of two unsupervised anomaly detection methods, ReContrast and RD++. These methods, without fine-tuning using labeled datasets, are referred to as “zero-shot” methods for convenience.
[…]
Diagnosis Performance on Retinal Fundus Images We further applied AFLoc to retinal fundus images, the results are presented in Table 2. The evaluation covered nine retinal diseases: Macular Degeneration (MD), Retinopathy, Myopia, Glaucoma, Congenital Optic Disc Anomalies (CODA), Retinal Arteriosclerosis (RAS), Cataract, Macular Epiretinal Membrane (MEM), and Macular Lesion (ML). While unsupervised anomaly detection methods, ReContrast and RD++, perform well on specific diseases like Myopia (0.921 AUROC) and Cataract (0.819 AUROC), their average scores are lower (0.589 and 0.52, respectively) compared to vision-language pre-training methods. Among the label free methods, AFLoc achieves the highest scores across all diseases, with an average AUROC of 0.908, much higher than the second-best method, GLoRIA (0.772). Additionally, AFLoc’s performance is comparable to RETFound, which used labeled data for fine-tuning. These results highlight AFLoc’s exceptional diagnostic capability and its potential for generalizing across diverse diseases.
[…]
a
b
| Methods | Label free | MD | Retinopathy | Myopia | Glaucoma | CODA | RAS | Cataract | MEM | ML | Mean |
| ReContrast | ✓ | 0.423 | 0.689 | 0.921 | 0.693 | 0.468 | 0.408 | 0.341 | 0.546 | 0.812 | 0.589 |
| RD++ | ✓ | 0.569 | 0.518 | 0.291 | 0.431 | 0.634 | 0.671 | 0.819 | 0.569 | 0.176 | 0.520 |
| ConVIRT | ✓ | 0.421 | 0.759 | 0.870 | 0.848 | 0.854 | 0.466 | 0.424 | 0.862 | 0.698 | 0.689 |
| GLoRIA | ✓ | 0.914 | 0.592 | 0.774 | 0.926 | 0.803 | 0.636 | 0.610 | 0.834 | 0.864 | 0.772 |
| AFLoc | ✓ | 0.978 | 0.839 | 0.962 | 0.938 | 0.899 | 0.643 | 0.987 | 0.978 | 0.946 | 0.908 |
| RETFound | ✗ | 0.935 | 0.758 | 0.971 | 0.918 | 0.900 | 0.901 | 0.995 | 0.908 | 0.883 | 0.908 |
Performance Enhancement via Limited Localization Annotations
This section aims to enhance AFLoc’s clinical applicability by presenting results obtained through leveraging limited labeled data for model fine-tuning on both segmentation and localization tasks.
As expected, AFLoc demonstrates enhanced segmentation performance with increased fine-tuning using labeled data on the SIIM dataset (Extended Data Table 9). The Dice score is increased from 0.772 with 1% of the data for fine-tuning to 0.809 with 100% of the data. Furthermore, AFLoc consistently achieves higher Dice scores than all comparative methods across varying portions of labeled fine-tuning data. These results indicate that AFLoc can effectively utilize limited annotations to achieve enhanced segmentation performance.
The results for pathology localization on the NIH ChestXray14 dataset are depicted in Extended Data Table 10. Consistent performance enhancement is achieved by AFLoc over the two comparative methods across different IoU thresholds. At an IoU of 0.3, AFLoc attains a mean localization accuracy of 0.62, surpassing the performance of the two comparative methods (Li et al. (0.49) and Wang et al. (0.22)). Notably, the model excels even under stricter thresholds like IoU = 0.5 and IoU = 0.7, especially for more challenging pathologies such as atelectasis and infiltration.
Extended Data Fig. 7 provides visual illustrations of AFLoc’s localization results on pneumothorax chest X-ray images and retinal hemorrhages. The model’s predictions align well with the ground truth, even for small and subtle lesions. These results demonstrate that AFLoc is capable of leveraging limited labeled data to improve its localization performance.
Overall, these findings suggest that AFLoc can further benefit from fine-tuning with annotated data, achieving competitive results in both segmentation and localization tasks. This capability is crucial for broadening AFLoc’s potential clinical applications.
Ablation Study of Text Granularities
In this section, we explore the impact of incorporating text features at different levels of granularity – word-level, sentence-level, and report-level – on the performance of AFLoc. The results are listed in Extended Data Table 11. When using only one level of text features, the performance of AFLoc deteriorates a lot, especially for word-level and report-level features. One possible explanation could be that the word-level features lack sufficient contextual information to capture the characteristics of the pathology, whereas the report-level features may fail to capture critical diagnostic details due to their overly macroscopic perspective. On the other hand, AFLoc using sentence-level features achieves relatively higher performance, suggesting that sentence-level features can provide richer contextual information, which is important for pathology localization in medical imaging. When these different levels of text features are combined in utilization, clear performance enhancement is observed, and the best performance is achieved when all three levels are employed, fully supporting our hypothesis that the combination of multi-level information is crucial for improving the accuracy of pathology localization in medical images.
Ablation Study of Text Encoders
Text encoders play a crucial role in deeply understanding medical concepts and achieving accurate semantic alignment. To investigate the impact of text encoding on pathology localization performance, we implemented the text encoder from the latest large language and vision assistant for biomedicine (LLaVA-Med [46]). LLaVA-Med has shown improvements in open biomedical question-answering tasks. However, as demonstrated in Extended Data Table 12, its performance falls short compared to BioClinicalBERT [47]. This discrepancy can be attributed to the superior performance of language models trained on domain-specific datasets (e.g., biomedical and clinical data) tailored for their respective tasks [48]. LLaVA-Med was trained on the PMC-15M dataset, which is extracted from scientific publications in PubMed Central, whereas BioClinicalBERT was trained on electronic health records of intensive care unit patients, closely aligning with the clinical reports used in this study. This domain-specific alignment likely provides BioClinicalBERT an advantage in comprehending intricate medical terminology and improving its localization performance.
Discussion
Pathology localization and diagnosis are essential components of clinical applications in medical imaging. However, existing methods encounter various challenges, notably stemming from the scarcity of annotated data related to generalization [19, 18]. In this study, we introduce AFLoc, an annotation-free deep learning model designed to enhance pathology localization by utilizing both clinical reports and imaging data in the pre-training phase.
AFLoc employs a multi-level semantic alignment strategy to enhance the medical feature representation. Given that pathologies often occupy only small regions within images, accurate localization can be challenging for global models. While some methods align global and local text-image features [30], they may miss critical semantic context by focusing excessively on individual words. In contrast, AFLoc adopts a more comprehensive approach by incorporating features at different levels of granularity (word-level, sentence-level, and report-level) during pre-training. This allows AFLoc to capture the necessary details for precise pathology localization.
We extensively evaluated AFLoc’s localization and diagnostic capabilities in an annotation-free setting across three major clinical imaging modalities: chest X-rays, retinal fundus, and histopathology. In the chest X-ray modality, AFLoc demonstrated improved localization performance compared to existing methods, such as MedKLIP and BioViL, across multiple evaluation metrics, including Intersection over Union (IoU), Dice similarity coefficient, and Contrast-to-Noise Ratio (CNR). AFLoc effectively localized a variety of thoracic pathologies, including pneumonia, atelectasis, and pleural effusion, on datasets such as RSNA Pneumonia, MS-CXR, and CheXlocalize. Notably, AFLoc showed good performance in localizing subtle and low-contrast pathologies, such as airspace opacity and atelectasis, which are often difficult to detect but crucial for early diagnosis and treatment planning. Furthermore, AFLoc surpassed the human benchmark on the CheXlocalize dataset across five key tasks, implying its potential to compete with conventional methods that depend on human-reviewed data. Additionally, AFLoc performed well on the COVID Rural dataset, which provides X-rays of patients affected with COVID-19 pneumonia, highlighting its potential for generalizing to new and emerging diseases. This capability is particularly important for timely and accurate disease detection, especially during global health crises like the COVID-19 pandemic.
The utilization of AFLoc with retinal fundus images underscores its versatility. In retinal pathology, AFLoc performed well at localizing and identifying features such as choroidal neovascularization and drusen, surpassing other methods like GLoRIA. Given the subtlety and diffuseness of retinal diseases, precise localization is particularly challenging. AFLoc’s adept performance without human-annotated data highlights its potential to improve ophthalmic diagnostics, especially in settings where large-scale data annotation is not feasible.
In histopathology, AFLoc demonstrated its effectiveness in detecting cancerous tissues and abnormal structures within prostate cancer samples, achieving better diagnosis performance than specialized models like PLIP and CONCH. Histopathological analysis is essential for cancer diagnosis and prognosis, as it involves the identification of cellular-level changes indicative of cancer. AFLoc’s success in this realm indicates its proficiency in handling intricate tissue morphology and capturing fine-grained details that are crucial for cancer detection and grading. This holds particular importance in clinical contexts, where pathologists rely on image-based assessments to make critical diagnostic decisions. AFLoc’s strong performance in this area highlights its potential to support pathologists by efficiently localizing pathological regions, allowing them to focus on higher-level diagnostic tasks.
Beyond its performance across different imaging modalities, AFLoc’s ability to outperform human benchmark performance on certain tasks (e.g., detecting opacity, atelectasis, edema, consolidation, and pleural effusion) further validates its potential for clinical deployment. These results suggest that AFLoc could serve as a complementary tool to human expertise, offering clinicians an additional support system that can help mitigate diagnostic errors and improve patient care. Moreover, AFLoc’s strong localization capabilities, even in complex or subtle cases, indicate its potential to automate some aspects of the diagnostic process, potentially enhancing efficiency in clinical practice.
Finally, we illustrated that AFLoc’s performance could be further improved with a small amount of annotated data, demonstrating its scalability and adaptability for clinical applications. While annotation-free localization is a valuable property, fine-tuning with labeled data can elevate AFLoc’s performance. This flexibility renders AFLoc fitting for diverse clinical environments, spanning well-resourced hospitals with access to large labeled datasets to smaller clinics in resource-limited settings. AFLoc’s adeptness in achieving commendable results with minimal annotated data makes it a potentially practical solution for healthcare systems globally, particularly in regions where access to expert annotators is limited.
In conclusion, AFLoc represents a promising step forward in medical image analysis by combining the strengths of multi-modal learning with annotation-free localization and diagnosis. Its strong performance across multiple modalities and its ability to assist in detecting both common and emerging diseases position it as a versatile tool with broad clinical applicability. With further refinements and integration into clinical workflows, AFLoc has the potential to improve medical diagnostics, providing healthcare professionals with AI-driven tools to enhance the accuracy and efficiency of patient care.
Despite these encouraging results, AFLoc does exhibit certain limitations. There is still room for improvement, particularly in refining more precise pathological segmentation. Implementing advanced image feature extraction techniques, like aligning local features of suggested regions more effectively with text features from multi-scale anchor boxes, may offer avenues to boost its accuracy. Additionally, AFLoc’s performance could benefit from access to larger, more diverse datasets. We view AFLoc as a potential benchmark for annotation-free methodologies in the medical field, with the possibility of playing a critical role in supporting healthcare systems during forthcoming disease outbreaks.
Methods
The Proposed AFLoc Model
Text encoding In clinical practice, medical texts may originate from various types of content, including clinical observations, diagnostic descriptions, and related information. These texts are typically composed of multiple sentences or phrases. In GLoRIA [30], a block tokenization technique was used, constructing complete words by aggregating multiple subwords. On the other hand, BioViL [28] employs a custom dictionary developed from multiple datasets, aiming to reduce the frequency of words being split into subwords. A radiological finding in an image might require information of location, size, and severity for a comprehensive description. Treating words as independent entities may overlook the complete semantics, leading to incorrect semantic alignment between text and image. In our proposed AFLoc, text features are extracted at three levels: word-level, sentence-level, and report-level, aiming to represent report content more comprehensively and precisely through multi-granularity semantics. Additionally, this three levels structure is naturally present in the text, requiring no complex designs to accomplish our text encoding.
Specifically, we employ BioClinicalBERT [47] as our text encoder to extract text features. When processing a medical report containing words and sentences, each word is tokenized into tokens. The tokenizer inputs the tokenized report into the text encoder , where H is the maximum token length. The text encoder outputs , L is the number of layers, and indicates the feature dimension. We take the average feature of the last 4 layers as the subword-level feature . Then, we aggregate features at the word, sentence, and report levels. Particularly, for a word, all its subword-level features are summarized to obtain the word-level features . On the other hand, all corresponding subword-level features are averaged to obtain sentence-level features and one report-level feature .
[…]
Multi-level semantic alignment The medical report is the textual description of the corresponding medical image, containing rich information at various levels of granularity to assist in diagnosis. Ideally, the semantic information in the medical image and its corresponding report should be consistent across these different levels. This motivates us to leverage multi-level semantic information to enhance cross-modal alignment. Specifically, AFLoc learns to align the shallow local features and deep local features from the image with word-level text features and sentence-level text features . In addition to local alignment, we also maintain global semantic alignment between the image’s global feature and report-level text features . This multi-level semantic alignment scheme is illustrated in Fig. 1.
Taking the alignment between sentence feature and deep local feature as an example. The similarity matrix between and can be computed by:
With the similarity , the attention weighted image feature can be represented based on the normalized similarity for a sentence across all deep local features:
Then the localized feature matching function can be represented as:
and are temperature parameters, and represents the size of the first dimension of the local features (i.e., for sentence-level features and for word-level features). is used to calculate the cosine similarity between the two vectors . The local contrastive loss between deep local features and sentence-level features is defined as:
where is a temperature parameter and is the batch size. Similarly, we obtain the contrastive loss between shallow local features and word-level features:
For global semantic alignment, we optimize the global features and report-level features according to the following loss function:
AFLoc is trained to jointly optimize the local and global semantic alignment. The final loss function of AFLoc is:
Annotation-free Pathology Localization and Diagnosis Once the model is trained, it can be applied to the pathology localization and clinical diagnosis tasks without requiring additional annotations. Instead, we transform the task into a text-image matching problem using text descriptions, offering greater flexibility as the text descriptions can be easily generated to meet various clinical requirements. By default, we use the automatically generated descriptions based on the pathology’s name and a template for testing. That is, for example, ‘Findings suggesting {lesion}’ for a positive description and ‘No findings suggesting {lesion}’ for a negative description. We refer to such description as ”Simple description”. In addition, AFLoc also supports the ”Precise description” of the pathology, such as ‘Large right-sided pneumothorax, air space opacity in a right infrahilar location’. This ”precise description” comes from clinicians who, based on their observation of the images, write corresponding descriptions to validate their hypotheses when the diagnosis is uncertain. With the text descriptions, the text features can be obtained with the pre-trained text encoder.
For the diagnosis task, we calculate the similarity between the global image features and the global text features for both positive and negative text descriptions. The positive similarity is then normalized to the range [0,1] as the prediction probability by softmax along the positive/negative dimension. The prediction will be considered as a positive prediction when the probability value is above a certain threshold.
For the localization task, we compute a similarity map between the deep local image features and the text features since it requires a more detailed representation. A bilinear interpolation and a normalization operation were applied to upscale the similarity map to the original image size and normalize it to the range of [-1, 1] as the generated heatmap for pathology localization. Finally, we use hard thresholding to obtain the binarized segmentation masks from the heatmaps. To ensure the reliability of the localization result, we only output the localization results with positive diagnosis predictions.
Experimental details We pre-train AFLoc on MIMIC-CXR, Retinal Fundus, and Quilt-1M dataset for chest X-ray, retinal fundus, and histopathology respectively. All images were uniformly resized to pixels and we followed the same image preprocessing strategy as [30].
A random shuffling and a sentence sampling strategy were applied to enhance the text data. Notably, the reports for the chest X-ray pretrain include multiple sections with different contents. We use the texts within the ’Findings’ and ’Impressions’ sections to train our model. Furthermore, we applied an additional sentence augmentation strategy to enhance the diversity of medical reports for retinal fundus data. Specifically, we generate five semantically equivalent but differently expressed sentences for each sentence by ChatGPT. These equivalent sentences will be randomly replaced during pre-training. For the histopathology dataset, we directly use the corresponding caption provided by the dataset for pre-training.
During pretraining, we used the Adam optimizer with an initial learning rate of 0.00002. A linear learning rate decay mechanism is employed by a decay factor of 0.9. The other hyperparameters were optimized with a corresponding validation dataset.
Evaluation metrics For the localization task, the Intersection over Union (IoU) and Dice coefficients were used to evaluate the thresholded results. We report the average results at multiple thresholds following the settings from BioViL. For the CheXlocalize dataset, we followed the methodology from [14] that determined an optimal threshold from the validation set and applied it to the test set. Additionally, we used the CNR [28] for measuring the difference in heatmap scores between inside and outside the bounding box area, eliminating the need for a hard threshold. This assessment of local similarity is significant, as certain clinical downstream applications might benefit more from heatmap visualizations than discrete segmentation masks. To directly assess localization performance without the influence of classification errors, we reported localization performance on the positive slices of the dataset (those with ground truth segmentations or bounding boxes annotated by radiologists). This approach ensured that the evaluation focused on the accuracy of localization, without being confounded by classification errors [14]. The averages of these metrics are reported on the test set, based on over 1,000 bootstrap repetitions, accompanied by the 95% confidence interval (CI), which was determined using the 2.5th and 97.5th percentiles of the empirical distribution.
For the diagnosis task, we reported the area under the receiver operating characteristic curve (AUROC) and balanced accuracy. The AUROC is calculated in the binary setting to reflect the overall classification performance at different thresholds. For multi-class classification, we computed the AUROC for each disease category and then averaged them to obtain the general AUROC. Balanced accuracy is defined as the macro average of recall for each class, which accounts for the class imbalance by treating each class equally.
Dataset
In our chest X-ray experiments, we utilized the MIMIC-CXR dataset for training. We then evaluated the model across six external datasets: RSNA, MS-CXR, CheXlocalize, COVID Rural, SIIM-ACR Pneumothorax, and ChestX-Det10. Additionally, to validate the model’s adaptability to different modalities, we collected retinal fundus data. For the retinal fundus experiments, the model was retrained using image-report pairs from this dataset.
We evaluated our model across three different imaging modalities: chest X-ray, retinal fundus, and histopathology. For the chest X-ray experiments, we pre-trained our model on the MIMIC-CXR dataset, then evaluated it across seven widely used external datasets: RSNA, MS-CXR, CheXlocalize, COVID Rural, SIIM-ACR Pneumothorax, NIH ChestXray14, and ChestX-Det10. For the retinal fundus experiments, we pre-trained and evaluated the model using a real-world in-house dataset. For the histopathology experiments, we pre-trained our model on the Quilt-1M dataset and then evaluated it on three external datasets: SICAPv2, WSSS4LUAD, and DHMC LUAD. The details of each dataset are provided below.
MIMIC-CXR [50]. The MIMIC-CXR dataset is an extensive, publicly available collection of chest X-ray images, paired with free-text radiological reports. It contains 377,110 images that are linked to 227,835 radiological studies. Each study within the database may consist of one or more images, providing a diverse range of scanning perspectives for analysis and research purposes.
RSNA Pneumonia [35]. The RSNA Pneumonia dataset presents a challenge in identifying visual signals of pneumonia from medical images. Pneumonia typically presents as one or multiple regions of increased opacity in chest X-rays [51]. However, diagnosing pneumonia based on chest X-rays can be complicated due to the presence of many other lung conditions, such as pulmonary edema, bleeding, atelectasis, lung cancer, or surgical changes. Additionally, pleural effusion outside the lungs can also appear as increased opacity in CXR. In this context, developing automated methods for pneumonia localization can be highly beneficial in the clinical environment. Images in this dataset are accompanied by bounding box labels that indicate the areas affected by pneumonia. Over 1000 cases in this dataset are positive, with each image potentially containing one or more boundary boxes highlighting the pneumonia regions.
MS-CXR [28]. MS-CXR is a curated database that features 1,153 image-sentence pairs of bounding boxes along with their corresponding phrases. This collection encompasses 8 different types of cardiopulmonary radiological findings. The image-sentence pairs are distributed equally among the 8 findings. In addition to brief descriptions, this dataset provides authentic depictions of radiological findings as described in radiology reports and oral records. All the data have been annotated and validated by professional radiologists, ensuring a gold standard for phrase grounding assessment. MS-CXR is a subset of the MIMIC-CXR dataset. All data in MS-CXR have been removed from the MIMIC-CXR training set to ensure it was not used in the training process.
CheXlocalize [14]. The CheXlocalize dataset is a comprehensively curated chest X-ray dataset. It features multi-label, pixel-level segmentations meticulously drawn by two board-certified radiologists, ensuring a highly precise ground truth annotation for each image. This dataset includes 234 images, encompassing 643 segmented pathologies. It covers a broad range of cardiopulmonary conditions, including airspace opacity, atelectasis, cardiomegaly, consolidation, edema, enlarged cardiomediastinum, lung lesion, pleural effusion, pneumothorax, and support devices. Moreover, CheXlocalize also provides human benchmark segmentations, drawn by a separate group of 3 board-certified radiologists, which can be used to study human benchmark localization performance.
COVID Rural [36, 37]. The COVID Rural project includes 221 chest X-ray images from 105 patients who tested positive for COVID-19. In these images, opaque areas associated with COVID-19 infection are delineated by polygons that connect specific points. The manual annotations of these images were generated by primary physicians and reviewed by experienced radiologists.
SIIM-ACR Pneumothorax (SIIM) [52]. The SIIM dataset is a widely adopted dataset for the evaluation of chest X-ray models. This dataset is designed to promote the development of models for identifying pneumothorax. In severe cases, pneumothorax can cause obstructive shock and be fatal if untreated [53]. This dataset includes more than 120,000 frontal chest X-rays, each of which has been precisely annotated. Employing the data partitioning method described by [19], we conducted tests on the test set.
NIH ChestXray14 [54]. This dataset contains 112,120 chest X-ray images, covering 14 different potentially concurrent chest diseases. For the classification task, we use the official data split for evaluation. For the localization task, we evaluate all 8 categories with localization annotations and follow the settings of [55] to ensure a fair comparison.
ChestX-Det10 [56]. This dataset is a subset of the NIH ChestXray14 dataset, consisting of 3,543 chest X-ray images. These images have been annotated with bounding boxes for 10 diseases/abnormalities by three board-certified radiologists. In this study, we followed the official data splits and conducted qualitative visual evaluations of our model on the test set.
Retinal Fundus. This dataset was compiled from a total of 26,028 retinal fundus images and paired reports from 11,372 individuals. All fundus images were captured using a Zeiss fundus camera and annotated by retinal experts with 10 years of experience for nine common retinal diseases. We divided the data set into training, validation, and testing sets by 8:1:1 based on the patient’s unique ID, which ensures that the same patient does not appear in the same data set. Our model was trained and validated on the training and validation sets, respectively, and classification performance evaluations were done on the testing dataset. To evaluate the performance of our model in pathology localization, we utilized an additional dataset for validation, which was annotated at the pixel level by retinal specialists. Specifically, this dataset includes 437 images of Choroidal neovascularization, 160 images of drusen, and 498 images of retinal hemorrhage.
Quilt-1M [41]. This dataset contains 1 million image-text pairs for histopathology. These data were collected from diverse sources, including medical Twitter, YouTube, and other internet repositories. The dataset comprises over 437,000 unique images paired with 802,000 corresponding text annotations, encompassing a total of 1.469 million UMLS entities extracted from the text. The images cover a range of microscopic magnification scales (0–10x, 10–20x, 20–40x) and were uniformly resized to 299×299 pixels for pretraining.
SICAPv2 [40]. The SICAPv2 dataset containing prostate histology whole slide images with both annotations of global Gleason scores and patch-level Gleason grades. We use the official test set with 2122 images from 31 whole slide images for both annotation-free localization and diagnosis tasks.
WSSS4LUAD [57, 58]. This dataset consists of 67 hematoxylin and eosin stained whole-slide images from Guangdong Provincial People’ Hospital (GDPH) and 20 whole slide images from the cancer genome atlas (TCGA). We filtered out 4693 image patches from the official training split following [45] for our evaluation. Each patch was labeled as tumor epithelial tissue, tumor-associated stroma tissue and normal tissue.
DHMC LUAD [59]. This dataset consists of 143 hematoxylin and eosin stained whole-slide images of lung adenocarcinoma from the Department of Pathology and Laboratory Medicine at Dartmouth-Hitchcock Medical Center (DHMC). All images were labeled into 5 categories, including lepidic, acinar, papillary, micropapillary, and solid. We use the entire dataset for annotation-free diagnosis task following [45].
Approvals for the research were secured from the Institutional Review Board/Ethics Committees of all participating hospitals or institutes. Consent forms were signed by all involved patients. The study was carried out in accordance with the United States Health Insurance Portability and Accountability Act. Furthermore, it was in compliance with the principles of the Declaration of Helsinki, and conformed to policies of the Chinese Center for Disease Control and Prevention, as well as the Chinese Health Laws.
Comparative Methods
In this section, we introduced the comparative algorithms we employed, which are categorized into four main groups: vision-language pre-training methods, self-supervised pre-training methods, unsupervised anomaly detection methods, and saliency-based methods. Each category will was discussed in detail in the following subsections.
Medical vision-language pre-training Methods Ten medical vision-language pre-training methods were compared in this study, including ConVIRT [60], GLoRIA [30], BioViL [28], and MedKLIP [19], CLIP [42], PLIP [44], BiomedCLIP [43], QuiltNet [41], CONCH [45]. ConVIRT improves visual representations by maximizing the agreement between true image-text pairs versus random pairs via a bidirectional contrastive objective between the image and text modalities. GLoRIA proposes an attention-based framework that learns global and local representations by comparing subregions of images with words in paired reports. BioViL achieves better text modeling by adopting an improved lexicon and innovative language pre-training methods. Then, contrastive learning is utilized with paired text and images. MedKLIP first extracts medical information from radiology reports through a triplet extraction module. Then, it uses a Transformer-based fusion model to combine this information and visual signals at the image block level, enhancing the capability of medical diagnostics. Other models are compared as state-of-the-art algorithms for histopathology, which can be found on Hugging Face: CLIP (https://huggingface.co/openai/clip-vit-base-patch16), PLIP (https://huggingface.co/vinid/plip), BiomedCLIP (https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224), QuiltNet (https://huggingface.co/wisdomik/QuiltNet-B-32), and CONCH (https://huggingface.co/MahmoodLab/CONCH). These models represent the latest advancements in medical vision-language pre-training. By leveraging large-scale datasets and cutting-edge pre-training techniques, they are capable of more accurate and efficient interpretation of medical images and associated texts, ultimately enhancing diagnostic capabilities.
Self-supervised Pre-training Methods RETFound [13] is a self-supervised pre-training model for retinal images that learns generalizable representations from unlabeled retinal images, providing a basis for label-efficient model adaptation across various applications. In this study, we utilized this model for transfer learning on the retinal fundus dataset using classification annotations and compared it with our proposed annotations-free AFLoc.
Unsupervised Anomaly Detection Methods This study compares ReContrast [24] and RD++ [23], two state-of-the-art unsupervised anomaly detection methods. ReContrast proposes a novel contrastive learning paradigm specifically designed for domain-specific unsupervised anomaly detection. This method enhances the transferability of the pre-trained encoder by jointly optimizing all parameters in an end-to-end fashion. The RD++ approach improves anomaly detection by increasing feature compactness and suppressing anomalous signals through a multi-task learning design. For the feature compactness task, RD++ introduces a self-supervised optimal transport method.
Saliency-based Methods In addition to medical vision-language pre-training methods, 3 comparative saliency methods were also adopted in our study. Grad-CAM [15] uses the gradients flowing into the final convolutional layer for any target category to generate a coarse localization map, highlighting the important areas in the image used for classification. Grad-CAM++ [16] is an extension of Grad-CAM that aims to enhance and refine the visualization of the decision-making process of convolutional neural networks. Compared to the original Grad-CAM approach, Grad-CAM++ can provide clearer and more detailed visual explanations, especially in identifying specific classes in recognition models. Eigen-CAM [17] is a visualization technique to interpret the decisions of convolutional neural networks. It extends the traditional Class Activation Mapping (CAM) method, aiming to offer a more intuitive and detailed visual interpretation. The core idea of Eigen-CAM is to use principal component analysis to extract significant features from the outputs of the convolutional layers.
Disease Localization Methods To evaluate the potential performance improvement of our method with limited location annotations, we compared with the methods proposed by Wang et al. [54] and Li et al. [55]. Wang et al. [54] proposed a weakly-supervised framework for weakly-supervised obect localization by considering large image capacity, various multi-label CNN losses and different pooling strategies. Li et al. [55] proposed a unified approach that effectively leverage class information and limited location annotations.
Acknowledgments
This research was partly supported by the National Natural Science Foundation of China (62222118, U22A2040), Shenzhen Medical Research Fund (B2402047), Shenzhen Science and Technology Program (RCYX20210706092104034, JCYJ20220531100213029), the major key project of Pengcheng Laboratory under grant PCL2023AS1-2, and Key Laboratory for Magnetic Resonance and Multimodality Imaging of Guangdong Province (2023B1212060052).
Data availability
MIMIC-CXR data is available at https://physionet.org/content/mimic-cxr/2.0.0 for users with credentialed access. RSNA Pneumonia data is available at https://www.rsna.org/rsnai/ai-image-challenge/RSNA-Pneumonia-Detection-Challenge-2018. MS-CXR data is available at https://physionet.org/content/ms-cxr/0.1/. CheXlocalize data is available at https://stanfordaimi.azurewebsites.net/datasets/abfb76e5-70d5-4315-badc-c94dd82e3d6d. COVID Rural data is available at https://www.cancerimagingarchive.net/collection/covid-19-ar/. SIIM-ACR Pneumothorax data is available at https://www.kaggle.com/c/siim-acr-pneumothorax-segmentation. NIH ChestXray4 data is available at https://nihcc.app.box.com/v/ChestXray-NIHCC. ChestX-Det10 data is available at https://github.com/Deepwise-AILab/ChestX-Det10-Dataset. Quilt-1M data is available at https://github.com/wisdomikezogwo/quilt1m. SICAPv2 data is available at https://data.mendeley.com/datasets/9xxm58dvs3/2. WSSS4LUAD data is available at https://wsss4luad.grand-challenge.org/. DHMC LUAD data is available at https://bmirds.github.io/LungCancer/.
Code availability
The code is available on GitHub at https://github.com/YH0517/AFLoc.
Appendix A
a
b
| Methods | RSNA Pneumonia | MS-CXR | Covid-Rural |
| IoU | |||
| ReContrast | 0.054 (0.050, 0.058) | 0.087 (0.072, 0.102) | 0.027 (0.020, 0.035) |
| RD++ | 0.050 (0.047, 0.054) | 0.090 (0.075, 0.105) | 0.058 (0.046, 0.072) |
| GLoRIA | 0.278 (0.269, 0.286) | 0.268 (0.242, 0.295) | 0.138 (0.119, 0.159) |
| BioViL | 0.295 (0.286, 0.304) | 0.228 (0.201, 0.255) | 0.191 (0.170, 0.215) |
| MedKLIP | 0.317 (0.309, 0.325) | 0.264 (0.240, 0.288) | 0.174 (0.156, 0.192) |
| AFLoc | 0.342 (0.332, 0.350) | 0.324 (0.298, 0.350) | 0.211 (0.185, 0.236) |
| Dice | |||
| ReContrast | 0.095 (0.089, 0.10) | 0.148 (0.125, 0.171) | 0.049 (0.037, 0.06) |
| RD++ | 0.087 (0.081, 0.093) | 0.152 (0.130, 0.176) | 0.100 (0.081, 0.119) |
| GLoRIA | 0.411 (0.400, 0.421) | 0.392 (0.357, 0.427) | 0.223 (0.193, 0.251) |
| BioViL | 0.433 (0.423, 0.443) | 0.342 (0.308, 0.377) | 0.298 (0.268, 0.329) |
| MedKLIP | 0.465 (0.455, 0.474) | 0.394 (0.365, 0.423) | 0.281 (0.252, 0.308) |
| AFLoc | 0.484 (0.474, 0.495) | 0.462 (0.431, 0.492) | 0.323 (0.289, 0.355) |
| CNR | |||
| ReContrast | 0.087 (0.064, 0.109) | 0.465 (0.398, 0.530) | -0.244 (-0.325, -0.165) |
| RD++ | 0.225 (0.201, 0.248) | 0.539 (0.474, 0.602) | 0.289 (0.214, 0.361) |
| GLoRIA | 1.223 (1.176, 1.272) | 1.287 (1.149, 1.421) | 0.663 (0.560, 0.770) |
| BioViL | 1.300 (1.262, 1.339) | 1.083 (0.949, 1.219) | 0.975 (0.880, 1.065) |
| MedKLIP | 1.032 (1.014, 1.051) | 1.044 (0.959, 1.132) | 0.838 (0.772, 0.906) |
| AFLoc | 1.538 (1.496, 1.580) | 1.636 (1.501, 1.772) | 1.062 (0.929, 1.173) |
| Methods | Choroidal neovascularization | Drusen | Intraretinal hemorrhage | Mean |
| IoU | ||||
| ReContrast | 0.080 (0.068, 0.093) | 0.027 (0.019, 0.035) | 0.035 (0.031, 0.038) | 0.047 (0.039, 0.055) |
| RD++ | 0.121 (0.112, 0.131) | 0.062 (0.051, 0.077) | 0.051 (0.047, 0.054) | 0.078 (0.070, 0.087) |
| ConVIRT | 0.131 (0.119, 0.144) | 0.120 (0.099, 0.140) | 0.079 (0.074, 0.085) | 0.110 (0.097, 0.123) |
| GLoRIA | 0.143 (0.130, 0.156) | 0.107 (0.085, 0.128) | 0.075 (0.070, 0.080) | 0.108 (0.095, 0.121) |
| AFLoc | 0.321 (0.303, 0.337) | 0.147 (0.125, 0.170) | 0.097 (0.091, 0.102) | 0.188 (0.173, 0.203) |
| Dice | ||||
| ReContrast | 0.124 (0.107, 0.141) | 0.046 (0.034, 0.061) | 0.064 (0.058, 0.069) | 0.078 (0.066, 0.090) |
| RD++ | 0.194 (0.179, 0.209) | 0.105 (0.087, 0.124) | 0.092 (0.087, 0.098) | 0.130 (0.118, 0.144) |
| ConVIRT | 0.205 (0.187, 0.223) | 0.187 (0.156, 0.215) | 0.141 (0.132, 0.149) | 0.178 (0.158, 0.199) |
| GLoRIA | 0.224 (0.205, 0.242) | 0.168 (0.137, 0.197) | 0.134 (0.126, 0.142) | 0.175 (0.156, 0.194) |
| AFLoc | 0.451 (0.430, 0.470) | 0.225 (0.192, 0.257) | 0.170 (0.160, 0.178) | 0.282 (0.261, 0.302) |
| CNR | ||||
| ReContrast | 0.901 (0.850, 0.954) | 0.421 (0.329, 0.510) | 0.253 (0.231, 0.274) | 0.525 (0.470, 0.579) |
| RD++ | 0.673 (0.635, 0.707) | 0.457 (0.395, 0.521) | 0.210 (0.191, 0.229) | 0.447 (0.407, 0.486) |
| ConVIRT | 0.359 (0.278, 0.447) | 0.652 (0.517, 0.778) | 0.255 (0.231, 0.278) | 0.422 (0.342, 0.501) |
| GLoRIA | 0.493 (0.427, 0.563) | 0.655 (0.512, 0.787) | 0.216 (0.194, 0.237) | 0.455 (0.378, 0.529) |
| AFLoc | 1.452 (1.368, 1.534) | 0.925 (0.783, 1.063) | 0.372 (0.349, 0.396) | 0.916 (0.833, 0.998) |
| Method | IoU | Dice | CNR |
| CLIP | 0.154 (0.141, 0.166) | 0.235 (0.218, 0.253) | 0.025 (-0.024, 0.074) |
| BiomedCLIP | 0.188 (0.173, 0.203) | 0.279 (0.260, 0.299) | 0.078 (0.004, 0.153) |
| PLIP | 0.249 (0.230, 0.269) | 0.352 (0.327, 0.376) | 0.321 (0.249, 0.398) |
| QuiltNet | 0.162 (0.147, 0.177) | 0.246 (0.226, 0.265) | 0.131 (0.073, 0.188) |
| CONCH | 0.198 (0.187, 0.210) | 0.300 (0.285, 0.316) | 0.378 (0.329, 0.426) |
| AFLoc | 0.285 (0.267, 0.304) | 0.399 (0.378, 0.421) | 0.630 (0.561, 0.697) |
| Disease | ReContrast | RD++ | GLoRIA | BioViL | MedKLIP | AFLoc |
| Pneumonia | 0.090 (0.080, 0.100) | 0.093 (0.084, 0.103) | 0.290 (0.265, 0.314) | 0.305 (0.284, 0.328) | 0.296 (0.277, 0.315) | 0.412 (0.386, 0.435) |
| Atelectasis | 0.079 (0.064, 0.095) | 0.075 (0.063, 0.088) | 0.303 (0.266, 0.339) | 0.266 (0.222, 0.312) | 0.324 (0.294, 0.355) | 0.381 (0.344, 0.417) |
| Consolidation | 0.098 (0.082, 0.115) | 0.120 (0.101, 0.139) | 0.304 (0.277, 0.330) | 0.299 (0.273, 0.325) | 0.265 (0.238, 0.289) | 0.371 (0.345, 0.395) |
| Pneumothorax | 0.053 (0.045, 0.062) | 0.093 (0.084, 0.103) | 0.116 (0.099, 0.133) | 0.109 (0.094, 0.124) | 0.092 (0.078, 0.106) | 0.117 (0.099, 0.135) |
| Pleural Effusion | 0.061 (0.048, 0.074) | 0.048 (0.037, 0.058) | 0.330 (0.304, 0.357) | 0.227 (0.205, 0.250) | 0.216 (0.195, 0.237) | 0.350 (0.326, 0.375) |
| Lung Opacity | 0.061 (0.046, 0.078) | 0.078 (0.060, 0.098) | 0.196 (0.164, 0.230) | 0.178 (0.148, 0.209) | 0.197 (0.167, 0.228) | 0.327 (0.290, 0.364) |
| Cardiomegaly | 0.131 (0.121, 0.140) | 0.073 (0.066, 0.080) | 0.408 (0.398, 0.420) | 0.267 (0.247, 0.286) | 0.394 (0.384, 0.404) | 0.367 (0.357, 0.377) |
| Edema | 0.120 (0.092, 0.150) | 0.162 (0.129, 0.193) | 0.201 (0.165, 0.236) | 0.171 (0.135, 0.206) | 0.326 (0.285, 0.366) | 0.269 (0.237, 0.299) |
| Dataset | Disease |
ReContrast |
RD++ |
ConVIRT |
GLoRIA |
BioViL |
CheXzero |
MedKLIP |
AFLoc |
| RSNA | Pneumonia | 0.632 | 0.566 | 0.804 | 0.715 | 0.828 | 0.858 | 0.869 | 0.881 (0.868, 0.893) |
| SIIM | Pneumothorax | 0.566 | 0.492 | 0.643 | 0.534 | 0.708 | 0.688 | 0.892 | 0.902 (0.890, 0.915) |
| NIH ChestX-ray14 | Mean | 0.511 | 0.486 | 0.560 | 0.610 | 0.662 | 0.679 | 0.726 | 0.737 (0.732, 0.741) |
| Atelectasis | 0.510 | 0.521 | 0.459 | 0.653 | 0.517 | 0.624 | 0.671 | 0.720 (0.710, 0.728) | |
| Cardiomegaly | 0.514 | 0.485 | 0.433 | 0.704 | 0.688 | 0.837 | 0.842 | 0.856 (0.845, 0.866) | |
| Effuion | 0.530 | 0.510 | 0.646 | 0.762 | 0.743 | 0.787 | 0.813 | 0.798 (0.791, 0.805) | |
| Infiltration | 0.553 | 0.502 | 0.654 | 0.660 | 0.601 | 0.651 | 0.706 | 0.684 (0.676, 0.691) | |
| Mass | 0.453 | 0.481 | 0.601 | 0.613 | 0.663 | 0.686 | 0.742 | 0.733 (0.720, 0.744) | |
| Nodule | 0.460 | 0.482 | 0.580 | 0.508 | 0.639 | 0.580 | 0.621 | 0.617 (0.602, 0.632) | |
| Pneumonia | 0.522 | 0.480 | 0.640 | 0.587 | 0.669 | 0.691 | 0.698 | 0.701 (0.681, 0.722) | |
| Pneumothorax | 0.535 | 0.445 | 0.533 | 0.572 | 0.683 | 0.793 | 0.821 | 0.843 (0.835, 0.851) | |
| Consolidation | 0.556 | 0.508 | 0.646 | 0.697 | 0.650 | 0.691 | 0.719 | 0.720 (0.708, 0.730) | |
| Edema | 0.541 | 0.542 | 0.692 | 0.762 | 0.795 | 0.804 | 0.803 | 0.810 (0.798, 0.822) | |
| Emphysema | 0.534 | 0.496 | 0.431 | 0.499 | 0.656 | 0.431 | 0.783 | 0.640 (0.627, 0.655) | |
| Fibrosis | 0.466 | 0.441 | 0.482 | 0.459 | 0.632 | 0.617 | 0.604 | 0.625 (0.598, 0.655) | |
| Pleural thickening | 0.491 | 0.499 | 0.545 | 0.613 | 0.637 | 0.524 | 0.499 | 0.718 (0.704, 0.733) | |
| Hernia | 0.485 | 0.417 | 0.494 | 0.450 | 0.698 | 0.783 | 0.841 | 0.847 (0.798, 0.884) |
| Method | SICAPv2 | WSSS4LUAD | DHMC LUAD |
| CLIP* | 0.250 (0.250, 0.256) | 0.333 (0.328, 0.333) | 0.194 (0.188, 0.205) |
| BiomedCLIP* | 0.375 (0.351, 0.389) | 0.512 (0.474, 0.550) | 0.253 (0.240, 0.273) |
| PLIP* | 0.287 (0.268, 0.324) | 0.462 (0.439, 0.492) | 0.231 (0.222, 0.246) |
| QuiltNet | 0.249 (0.226, 0.327) | 0.445 (0.429, 0.461) | 0.208 (0.089, 0.347) |
| CONCH* | 0.339 (0.325, 0.396) | 0.598 (0.560, 0.650) | 0.314 (0.291, 0.346) |
| AFLoc | 0.497 (0.471, 0.523) | 0.679 (0.663, 0.693) | 0.504 (0.371, 0.703) |
| Data Portion | ConVIRT | GLoRIA | BioViL | MedKLIP | AFLoc |
| 1% | 0.541 | 0.567 | 0.627 | 0.666 | 0.772 |
| 10% | 0.612 | 0.578 | 0.700 | 0.721 | 0.781 |
| 100% | 0.735 | 0.769 | 0.785 | 0.794 | 0.809 |
| T(IoU) | Method |
Atelectasis |
Cardiomegaly |
Effusion |
Infiltration |
Mass |
Nodule |
Pneumonia |
Pneumothorax |
Mean |
| 0.1 | Wang et al. [54] | 0.69 | 0.94 | 0.66 | 0.71 | 0.40 | 0.14 | 0.63 | 0.38 | 0.57 |
| Li et al. [55] | 0.71 | 0.98 | 0.87 | 0.92 | 0.71 | 0.40 | 0.60 | 0.63 | 0.73 | |
| AFLoc | 0.72 | 1.00 | 0.86 | 0.91 | 0.78 | 0.68 | 0.90 | 0.82 | 0.83 | |
| 0.2 | Wang et al. [54] | 0.47 | 0.68 | 0.45 | 0.48 | 0.26 | 0.05 | 0.35 | 0.23 | 0.37 |
| Li et al. [55] | 0.53 | 0.97 | 0.76 | 0.83 | 0.59 | 0.29 | 0.50 | 0.51 | 0.62 | |
| AFLoc | 0.59 | 1.00 | 0.69 | 0.86 | 0.66 | 0.57 | 0.78 | 0.65 | 0.73 | |
| 0.3 | Wang et al. [54] | 0.24 | 0.46 | 0.30 | 0.28 | 0.15 | 0.04 | 0.16 | 0.13 | 0.22 |
| Li et al. [55] | 0.36 | 0.94 | 0.56 | 0.66 | 0.45 | 0.17 | 0.39 | 0.44 | 0.49 | |
| AFLoc | 0.43 | 1.00 | 0.49 | 0.72 | 0.57 | 0.52 | 0.73 | 0.45 | 0.62 | |
| 0.4 | Wang et al. [54] | 0.09 | 0.28 | 0.20 | 0.12 | 0.07 | 0.01 | 0.08 | 0.07 | 0.12 |
| Li et al. [55] | 0.25 | 0.88 | 0.37 | 0.50 | 0.33 | 0.11 | 0.26 | 0.29 | 0.42 | |
| AFLoc | 0.30 | 1.00 | 0.26 | 0.52 | 0.46 | 0.34 | 0.57 | 0.31 | 0.47 | |
| 0.5 | Wang et al. [54] | 0.05 | 0.18 | 0.11 | 0.07 | 0.01 | 0.01 | 0.03 | 0.03 | 0.06 |
| Li et al. [55] | 0.14 | 0.84 | 0.22 | 0.30 | 0.22 | 0.07 | 0.17 | 0.19 | 0.27 | |
| AFLoc | 0.19 | 0.99 | 0.12 | 0.32 | 0.33 | 0.25 | 0.38 | 0.21 | 0.35 | |
| 0.6 | Wang et al. [54] | 0.02 | 0.08 | 0.05 | 0.02 | 0.00 | 0.01 | 0.02 | 0.03 | 0.03 |
| Li et al. [55] | 0.07 | 0.73 | 0.15 | 0.18 | 0.16 | 0.03 | 0.10 | 0.12 | 0.19 | |
| AFLoc | 0.09 | 0.94 | 0.05 | 0.21 | 0.22 | 0.19 | 0.20 | 0.11 | 0.25 | |
| 0.7 | Wang et al. [54] | 0.01 | 0.03 | 0.02 | 0.00 | 0.00 | 0.00 | 0.01 | 0.02 | 0.01 |
| Li et al. [55] | 0.04 | 0.52 | 0.07 | 0.09 | 0.11 | 0.01 | 0.05 | 0.05 | 0.12 | |
| AFLoc | 0.03 | 0.78 | 0.02 | 0.07 | 0.14 | 0.08 | 0.07 | 0.02 | 0.15 |
| Different Levels of Features | RSNA Pneumonia | COVID Rural | MS-CXR | CheXlocalize | Mean | ||
| Word – Shallow local | Sentence – Deep local | Report – Global | |||||
| ✓ | 0.239 | 0.147 | 0.180 | 0.133 | 0.175 | ||
| ✓ | 0.281 | 0.169 | 0.318 | 0.301 | 0.267 | ||
| ✓ | 0.191 | 0.056 | 0.105 | 0.186 | 0.135 | ||
| ✓ | ✓ | 0.329 | 0.177 | 0.264 | 0.285 | 0.264 | |
| ✓ | ✓ | 0.254 | 0.159 | 0.175 | 0.200 | 0.197 | |
| ✓ | ✓ | 0.285 | 0.211 | 0.325 | 0.305 | 0.282 | |
| ✓ | ✓ | ✓ | 0.342 | 0.212 | 0.324 | 0.318 | 0.299 |
| Text encoder | IoU | Dice | CNR |
| LLaVA-Med-v1.5-Mistral-7B | 0.307 | 0.441 | 1.483 |
| BioClinicalBERT | 0.324 | 0.461 | 1.634 |